# QCPINN: Quantum-Classical Physics-Informed Neural Networks for Solving PDEs

### v5

@@ Abstract
Physics-informed neural networks (PINNs) have emerged as promising methods for solving partial differential equations (PDEs) by embedding physical laws within neural architectures. However, these classical approaches often require a large number of parameters to achieve reasonable accuracy, particularly for complex PDEs. In this paper, we present a quantum-classical physics-informed neural network (QCPINN) that combines quantum and classical components, allowing us to solve PDEs with significantly fewer parameters while maintaining comparable accuracy and convergence to classical PINNs. We systematically evaluated two quantum circuit architectures across various configurations on five benchmark PDEs to identify optimal QCPINN designs. Our results demonstrate that the QCPINN achieves stable convergence and comparable accuracy, while requiring approximately 10% of the trainable parameters used in classical approaches. It also results in a 40% reduction in the relative error $L_{2}$ for the convection-diffusion equation. These findings demonstrate the potential of parameter efficiency as a measurable quantum advantage in physics-informed machine learning, significantly reducing model complexity while preserving solution quality. This approach presents a promising solution to the computational challenges associated with solving PDEs.
@@ 1 Introduction
Recent advancements in machine learning have transformed approaches to solving PDEs, with PINNs emerging as an innovative technique. PINNs embed physical laws directly into neural network architectures, eliminating the need for labeled training data while strictly enforcing physical constraints. This approach has demonstrated considerable success in modeling various physical phenomena across scientific and engineering domains [1, 2, 3]. Although they hold great potential, research is underway to improve the accuracy of PINNs in modeling different PDEs and physical systems. They often face challenges with stiff systems, capturing steep gradients, and efficiently navigating high-dimensional parameter spaces. While increasing model expressivity through additional parameters can alleviate these challenges, this approach introduces new problems: increased computational cost, risk of overfitting, and unstable convergence behavior during training [4, 5]. Such a trade-off between model complexity, generalization, and convergence ability is a fundamental challenge in neural network design.
@@
Quantum computing presents a promising paradigm for overcoming these limitations. With its ability to process high-dimensional data and exploit quantum principles such as superposition and entanglement, quantum algorithms offer unique advantages for solving complex computational problems. For instance, the Harrow-Hassidim-Lloyd (HHL) algorithm has demonstrated theoretical exponential speedups for specific linear systems, while variational quantum algorithms provide practical near-term solutions [6, 7, 8]. Moreover, recent developments [9, 10, 11] indicate that quantum neural networks possess an expressive power that could significantly enhance the learning capabilities for modeling complex physical systems.
@@
Although both PINNs and quantum computing have advanced independently, their integration remains largely unexplored. Several preliminary efforts have emerged: Sedykh et al. proposed HQPINN [12], focusing on computational fluid dynamics in 3D Y-shaped mixers; Dehaghani et al. introduced QPINN [13], integrating a dynamic quantum circuit with classical computing methodologies for optimal control problems; Trahan et al. [14] investigated both purely quantum and hybrid PINNs, highlighting parameter efficiency but limiting application to parametrized quantum circuit (PQC) networks with simple PDEs. Leong et al. proposed HQPINN [15], which is inspired by Fourier neural operators. However, existing research has not comprehensively established whether quantum-enhanced models can improve upon purely classical PINNs across a broad range of PDEs efficiently and reliably.
@@
We aim to investigate the limitations of traditional PINNs by integrating quantum computing circuits with classical neural networks. We propose that such an integrated or hybrid approach can leverage complementary strengths: classical computing’s robustness and quantum computing’s enhanced expressivity and inherent parallelism. We believe this method will allow us to solve a wide range of PDEs with similar or better accuracy and convergence rates compared to classical PINNs, while requiring significantly fewer trainable parameters than classical models.
@@
To this end, we introduce a hybrid architecture named Quantum-Classical Physics-Informed Neural Network (QCPINN). To search for a suitable QCPINN architecture and its configurations, we systematically evaluate multiple dimensions of the design space: two quantum circuit architectures: (a) discrete-variable (DV) qubit-based circuits with configurations, such as four circuit topologies: alternate, cascade, cross-mesh, and layered; and two embedding schemes: amplitude and angle; (b) continuous-variable (CV) circuits with configurations, including two measurement schemes: number and quadrature; three nonlinear operations: Kerr, cubic phase, and cross-Kerr; and two parameterizations: phase-free (magnitude) and full (magnitude and phase). We evaluate the performance of these QCPINN models across various PDEs, such as the Helmholtz and lid-driven cavity equations, while also extending our analysis to the Klein-Gordon, wave, and convection-diffusion equations. Our comprehensive assessment compares (a) solution accuracy, (b) convergence rates, and (c) parameter efficiency between classical PINNs and QCPINN variants.
@@
To ensure reproducibility and facilitate broader adoption of our methods, we have made all components of QCPINN open-source code and publicly available on GitHub at https://github.com/afrah/QCPINN. We hope this repository will enable community engagement and collaborative advancement of this emerging field.
@@ 2.1 Physics Informed Neural Network (PINNs)
PINNs [16] provide a framework to embed governing physical laws directly into neural network architectures. The PINN loss function combines terms representing governing equations and boundary conditions designed to minimize deviations from physical laws. More formally, the PINN loss function can be defined as:
$\begin{split}\mathcal{L}(\theta)=&\arg\min_{\theta}\sum_{k=1}^{n}\lambda_{k}\mathcal{L}_{k}(\theta)\\ =&\arg\min_{\theta}\left(\lambda_{1}\mathcal{L}_{1}(D[u_{\theta}(\mathbf{x});\mathbf{a}]-f(\mathbf{x}))+\sum_{k=2}^{n}\lambda_{k}\mathcal{L}_{k}(B[u_{\theta}(\mathbf{x})]-g_{k}(\mathbf{x}))\right),\end{split}$ (1)
where, $n$ denotes the total number of loss terms, $\lambda_{k}$ are the weighting coefficients, and $\theta$ represents trainable parameters. Here, $D[u_{\theta}(\mathbf{x});\mathbf{a}]$ and $B[u_{\theta}(\mathbf{x})]$ are the arbitrary differential operators and initial/boundary conditions, respectively, that are applied to the output of the neural network $u_{\theta}(\mathbf{x})$. While the initial and boundary conditions ensure that the network output satisfies predefined constraints at specific spatial or temporal locations, the physics loss acts as a regularizer, guiding the network to learn physically consistent solutions and preventing overfitting to sparse or noisy data. The balance between these loss components is crucial for achieving accurate solutions, as improper weighting can lead to underfitting the initial/boundary conditions or failing to capture the system dynamics correctly.
@@ 2.2 Continuous Variable (CV) QNN
In photonic quantum computing, CV computation addresses quantum states with continuous degrees of freedom, such as the position ($x$) and momentum ($p$) quadratures of the electromagnetic field. The computational space is, in principle, unbounded, involving infinite-dimensional Hilbert spaces with continuous spectra. To make these simulations computationally feasible, a cutoff dimension is employed to limit the number of basis states used for approximation. The computational state space is defined as $r=n^{m}$, where $m$ represents the number of qumodes and $n$ is the cutoff dimension. For example, with three qumodes and a cutoff dimension of 2, the state space is 8-dimensional. Ideally, higher cutoff dimensions (up to a certain threshold) yield more precise results but at the cost of increased computational complexity.
@@
where, $M$ represents the combined symplectic matrix of the interferometers and squeezing gates, $x$ is the input vector in the phase space, $b$ is the displacement vector, and $\Phi$ is the nonlinear transformation.
In addition to the building blocks above, CV-circuit QNNs offer flexibility in measurement schemes and parameterization strategies. Measurement can be performed in either the number basis, which projects the state onto discrete photon-number states, or the quadrature basis, which yields continuous-valued outcomes corresponding to position or momentum observables. Furthermore, the parameterization of unitary operations may be either full (including both magnitude and phase) or restricted to magnitude-only (by setting phase parameters $\phi$ to zero), which can improve trainability and mitigate optimization challenges at the cost of reduced expressivity.
Although such a model exhibits intriguing theoretical properties, their practical implementation in real-world problem solving seems to suffer from fundamental stability issues. This is mainly due to their inherent sensitivity and reliance on continuous degrees of freedom, which complicate optimization compared to DV-based approaches (discussed in detail in section 2.3). This sensitivity stems from several factors, including the precision required for state preparation, noise from decoherence, and the accumulation of errors in variational quantum circuits. Furthermore, the transformations involved in CV quantum circuits are prone to numerical instabilities, particularly when optimized using gradient-based methods, as seen in the following section. Although foundational studies about the barren plateau problem, such as [18, 19, 20], primarily address DV circuit systems, the analysis indicates that similar issues can arise in CV quantum systems. In particular, the infinite-dimensional Hilbert space and the continuous nature of parameterization can exacerbate barren-plateau effects.
@@ 2.3 Discrete Variable (DV) QNN
DV-circuit quantum models do not inherently provide affine transformations as a core feature like CV-based QNNs. However, they can achieve affine-like behaviour through parameterized single-qubit rotations, controlled operations, and phase shifts for amplitude adjustments. Therefore, the affine transform of equation (2) can be reinterpreted to align with their architecture such that:
@@
    - $U_{1},U_{2}$: A unitary operator that can be represented by parameterized single-qubit rotations, such as $R_{x}(\theta),R_{y}(\theta)$, or $R_{z}(\theta)$.
    - $S$: A sequence of controlled gates (e.g., CNOT and CZ) responsible for introducing entanglement, which enables the representation of correlated variables and interactions.
@@
3 Related Work
3.1 Exclusively Quantum Algorithms
Despite these advancements, purely quantum algorithms for solving PDEs encounter practical challenges such as requirements for quantum error correction [34, 35, 36], limitations in circuit depth [37, 31, 38], and bottlenecks in data embedding [39, 40, 41]. These obstacles, among others, limit the near-term viability of exclusively quantum PDE solvers on today’s noisy intermediate-scale quantum (NISQ) devices.
3.2 Hybrid Quantum-Classical Approaches
Hybrid quantum-classical approaches have emerged as promising solutions to overcome the limitations of exclusively quantum algorithms. These methods combine the robustness, maturity, and stability of classical computation with the enhanced capacity, expressivity, and inherent parallelism of quantum systems. While classical components can efficiently handle pre-/post-processing, optimization, and error stabilization, quantum layers can capture intricate patterns, solve sub-problems, and accelerate specialized computations.
Within this hybrid paradigm, Quantum Neural Networks (QNNs) have demonstrated universal approximation capabilities for continuous functions when implemented with sufficient depth, width, and well-designed quantum feature maps [42, 43, 44]. For instance, Salinas et al. [45] showed that single-qubit networks with iterative input reuploading can approximate bounded complex functions, while Goto et al. [46] established universal approximation theorems for quantum feedforward networks. Schuld et al. [9] revealed that the outputs of parameterized quantum circuits (with suitable data embedding and circuit depth) can be expressed as truncated Fourier series, thereby providing theoretical support for the universal approximation ability of QNNs. Liu et al. [47] provided formal evidence of quantum advantage in function approximation. These developments not only support applications in solving PDEs and optimization tasks but also extend to other computational tasks such as quantum chemistry [48], material sciences [49, 50], and pattern recognition [51], to name a few.
A well-known example of this integrated approach is the Variational Quantum Algorithms (VQA), also known as Variational Quantum Circuits (VQC). These algorithms employ parameterized quantum circuits with tunable gates optimized through classical feedback loops. VQA implementations for solving PDEs include the Variational Quantum Eigensolver (VQE)[52, 53], Variational Quantum Linear Solver (VQLS)[54], and Quantum Approximate Optimization Algorithm (QAOA)[55], as well as emerging methods like Differentiable Quantum Circuits (DQCs)[56, 57], quantum kernel methods [58], and Quantum Nonlinear Processing Algorithms (QNPA) [59].
The aforementioned methods primarily utilize qubit-based quantum computing, which represents the predominant paradigm in quantum computing literature. However, continuous-variable quantum computing (CVQC) [60] offers an alternative computational model that leverages continuous degrees of freedom, such as quadratures and electromagnetic field momentum as discussed in section 2.2. For instance, in [61], Gaussian and non-Gaussian gates were used to manage continuous degrees of freedom, offering enhanced scalability and expressivity for continuous systems. For consistency, we refer to qubit-based models as discrete-variable (DV) circuit models throughout this paper.
Recently, VQA methods have been extended to Physics-informed quantum neural networks (PIQNN) [62, 13, 14, 12, 15], broadening their applicability across physics and engineering domains. These hybrid methods offer several distinct advantages when implemented within the PINN framework.
Firstly, these models enable the flexible configuration of inputs and outputs independent of qubit/qumode counts or quantum layer depth, making them particularly effective for scenarios where the input dimension (e.g., temporal and spatial coordinates) differs from output dimensions (e.g., dependent variables like velocity, pressure, and force).
Secondly, hybrid approaches facilitate increased trainable parameter counts without the need for deeper quantum circuits, additional qubits, or higher cutoff dimensions in CV quantum systems. These factors are particularly crucial in the NISQ era, where hardware constraints impose strict limits on circuit depth and coherence times, making deeper quantum networks computationally expensive and impractical [63, 64].
Thirdly, these qualities are especially advantageous for PINNs, where shallow network architectures are favored to optimize the balance between computational efficiency and solution accuracy [65, 66, 67, 5].
Finally, a substantial benefit of such hybrid models is the integration of automatic differentiation within both classical and quantum networks. This capability, supported by contemporary software packages like PennyLane [68] and TensorFlow Quantum [69], enables the automatic computation of derivatives as variables are directly incorporated into the computational graph. This approach simplifies PINN implementation and facilitates the efficient handling of complex physics-informed loss functions and higher-order derivatives.
Our work is related to the approaches in [12, 13, 14, 15], but with several key differences. While HQPINN [12] focuses specifically on computational fluid dynamics in 3D Y-shaped mixers using a fixed quantum depth with an infused layer structure, the QPINN [13] integrates a dynamic quantum circuit combining Gaussian and non-Gaussian gates with classical computing methodologies within a PINN framework to solve optimal control problems. In addition, the model proposed in [14] investigates both purely quantum and hybrid PINNs, highlighting parameter efficiency but relying on limited PQC networks and addressing primarily simple PDEs. Recently proposed HQPINN [15], inspired by Fourier neural operator (FNO) [70], introduces a parallel hybrid architecture to separately model harmonic and non-harmonic features before combining them for output.

### v6

@@ Abstract
Physics-informed neural networks (PINNs) have emerged as promising methods for solving partial differential equations (PDEs) by embedding physical laws within neural architectures. However, these classical approaches often require a large number of parameters to achieve reasonable accuracy, particularly for complex PDEs. In this paper, we present a quantum-classical physics-informed neural network (QCPINN) that combines quantum and classical components, allowing us to solve PDEs with significantly fewer parameters while maintaining comparable accuracy and convergence to classical PINNs. We systematically evaluated two quantum circuit architectures across various configurations on five benchmark PDEs to identify optimal QCPINN designs. Our results demonstrate that the QCPINN achieves stable convergence and comparable accuracy, while using only 10-30% of the trainable parameters required by classical PINNs. This approach also results in a significant reduction in the relative $L_{2}$ error for Helmholtz, Klein-Gordon, and Convection-diffusion equations, with reduction ranging from 4% to 64% across various fields. These findings demonstrate the potential of parameter efficiency and solution accuracy in physics-informed machine learning, allowing for a substantial decrease in model complexity without compromising solution quality. QCPINN presents a promising pathway to address the computational challenges associated with solving PDEs.
@@ 1 Introduction
Recent advancements in machine learning have transformed approaches to solving partial differential equations (PDEs), with Physics-informed neural networks (PINNs) emerging as an innovative technique. PINNs embed physical laws directly into neural network architectures, eliminating the need for labeled training data while strictly enforcing physical constraints. This approach has demonstrated considerable success in modeling various physical phenomena across scientific and engineering domains [1, 2, 3, 4]. Although they hold great potential, research is underway to improve the accuracy of PINNs in modeling different PDEs and physical systems. They often encounter challenges with stiff systems, capturing steep gradients, and efficiently navigating high-dimensional parameter spaces for reasons such as memory requirements, slow convergence, or when calculating high-order or high-dimensional derivatives [5, 6]. While increasing model expressivity through additional parameters can alleviate these challenges, this approach introduces new problems: increased computational cost, risk of overfitting, and unstable convergence behavior during training [7, 8]. Such a trade-off between model complexity, generalization, and convergence ability is a fundamental challenge in neural network design.
@@
Quantum computing presents a promising paradigm for overcoming these limitations. With its ability to process high-dimensional data and exploit quantum principles such as superposition and entanglement, quantum algorithms offer unique advantages for solving complex computational problems. For instance, the Harrow-Hassidim-Lloyd (HHL) algorithm has demonstrated theoretical exponential speedups for specific linear systems, while variational quantum algorithms provide practical near-term solutions [9, 10, 11]. Moreover, recent developments [12, 13, 14, 15] indicate that quantum neural networks possess an expressive power that could significantly enhance the learning capabilities for modeling complex physical systems.
@@
Although both PINNs and quantum computing have advanced independently, their integration remains largely unexplored. Several preliminary efforts have emerged: Sedykh et al. proposed HQPINN [16], focusing on computational fluid dynamics in 3D Y-shaped mixers; Dehaghani et al. introduced QPINN [17], integrating a dynamic quantum circuit with classical computing methodologies for optimal control problems; Trahan et al. [18] investigated both purely quantum and hybrid PINNs. Leong et al. proposed HQPINN [19], which is inspired by Fourier neural operators. However, existing research has not comprehensively established whether quantum-enhanced models can improve upon purely classical PINNs across a broad range of PDEs efficiently and reliably.
@@
We investigate the limitations of traditional PINNs by integrating quantum computing circuits with classical neural networks. We propose that such a hybrid approach can leverage complementary strengths: classical computing’s robustness and quantum computing’s enhanced expressivity with inherent parallelism. We believe this method will allow us to solve a wide range of PDEs with similar or better accuracy while requiring significantly fewer trainable parameters than classical models.
@@
To this end, we introduce a hybrid architecture named Quantum-Classical Physics-Informed Neural Network (QCPINN). To search for a suitable QCPINN architecture and its configurations, we systematically evaluate multiple dimensions of the design space: two quantum circuit architectures: (a) discrete-variable (DV) qubit-based circuits with configurations, such as four circuit topologies: Alternate, Cascade, Cross-mesh, and Layered; and two embedding schemes: Amplitude and Angle; (b) continuous-variable (CV) circuits with configurations, including two measurement schemes: number and quadrature; three nonlinear operations: Kerr, cubic phase, and cross-Kerr; and two parameterizations: phase-free (magnitude) and full (magnitude and phase). We evaluate the performance of these QCPINN models across various PDEs, such as the Helmholtz and lid-driven Cavity equations, while also extending our analysis to the Klein-Gordon, Wave, and Convection-diffusion equations. Our comprehensive assessment compares (a) solution accuracy, (b) convergence rates, (c) parameter efficiency (d) time, and (f) memory between classical PINNs and QCPINN variants.
@@
To ensure reproducibility and facilitate collaborative advancement, we have made all components of QCPINN open-source code and publicly available on GitHub at https://github.com/afrah/QCPINN [20].
2 Related Work
2.1 Exclusively Quantum Algorithms
Despite these advancements, purely quantum algorithms for solving PDEs encounter practical challenges such as requirements for quantum error correction [32, 33, 34], limitations in circuit depth [35, 29, 36], and bottlenecks in classical data embedding [37, 38, 39]. These obstacles, among others, limit the near-term viability of exclusively quantum PDE solvers on today’s noisy intermediate-scale quantum (NISQ) devices.
2.2 Hybrid Quantum-Classical Approaches
Hybrid quantum-classical approaches have emerged as promising solutions to overcome the limitations of exclusively quantum algorithms. These methods combine the nonlinearity, maturity, and stability of classical computation with the enhanced capacity, expressivity, and inherent parallelism of quantum systems. While classical components can efficiently handle pre-/post-processing, optimization, and error stabilization, quantum layers can capture intricate patterns, solve sub-problems, and accelerate specialized computations.
Within this hybrid paradigm, Quantum Neural Networks (QNNs), also referred to as parameterized quantum circuits (PQC), have demonstrated universal approximation capabilities for continuous functions when implemented with sufficient depth, width, and well-designed quantum feature maps. For instance, Salinas et al. [40] showed that single-qubit networks with iterative input reuploading can approximate bounded complex functions, while Goto et al. [41] established universal approximation theorems for quantum feedforward networks. Schuld et al. [13] revealed that the outputs of parameterized quantum circuits (with suitable data embedding and circuit depth) can be expressed as truncated Fourier series, thereby providing theoretical support for the universal approximation ability of QNNs. Liu et al. [42] provided formal evidence of quantum advantage in function approximation. Killoran et al [43] demonstrate the universality of the continuous variable quantum models by showing that many classical neural networks can be effectively emulated on a photonic quantum computer. These developments not only support applications in solving PDEs and optimization tasks but also extend to other computational tasks such as quantum chemistry [44], material sciences [45, 46], and pattern recognition [47], to name a few.
A well-known example of this integrated approach is the Variational Quantum Algorithms (VQA) or Variational Quantum Circuits (VQC). These algorithms employ parameterized quantum circuits with tunable gates optimized through classical feedback loops. VQA implementations for solving PDEs include the Variational Quantum Eigensolver (VQE) [48, 49], Variational Quantum Linear Solver (VQLS) [50], and Quantum Approximate Optimization Algorithm (QAOA) [51], as well as emerging methods like Differentiable Quantum Circuits (DQCs) [52, 53], quantum kernel methods [54], and Quantum Nonlinear Processing Algorithms (QNPA) [55].
The aforementioned methods primarily utilize qubit-based quantum computing, which represents the predominant paradigm in quantum computing literature. However, continuous-variable quantum computing (CVQC) [56] offers an alternative computational model that leverages continuous degrees of freedom, such as quadratures and electromagnetic field momentum. For consistency, we refer to qubit-based models as discrete-variable (DV) circuit models throughout this work.
Recently, VQA have been extended to Physics-informed quantum neural networks (PIQNN) [57, 17, 18, 16, 19], broadening their applicability across physics and engineering domains. These hybrid methods offer several distinct advantages when implemented within the PINN framework.
Firstly, these models enable the flexible configuration of inputs and outputs independent of qubit/qumode counts or quantum layer depth, making them particularly effective for scenarios where the input dimension (e.g., temporal and spatial coordinates) differs from output dimensions (e.g., dependent variables like velocity, pressure, and force). Secondly, hybrid approaches facilitate increased trainable parameter counts without the need for deeper quantum circuits, additional qubits, or higher cutoff dimensions in CV quantum systems. These factors are particularly crucial in the NISQ era, where hardware constraints impose strict limits on circuit depth and coherence times, making deeper quantum networks computationally expensive and impractical [58, 59]. Thirdly, these qualities are especially advantageous for PINNs, where shallow network architectures are favored to optimize the balance between computational efficiency and solution accuracy [60, 61, 62, 8]. Finally, a substantial benefit of such hybrid models is the integration of automatic differentiation within both classical and quantum networks. This capability, supported by contemporary software packages like PennyLane [63] and TensorFlow Quantum [64], enables the automatic computation of derivatives as variables are directly incorporated into the computational graph. This approach simplifies PINN implementation and facilitates the efficient handling of complex physics-informed loss functions and higher-order derivatives.
Our work is related to the approaches in [16, 17, 18, 19], but with several key differences. While HQPINN [16] focuses specifically on computational fluid dynamics in 3D Y-shaped mixers using a fixed quantum depth with an infused layer structure, the QPINN [17] integrates a dynamic quantum circuit combining Gaussian and non-Gaussian gates with classical computing methodologies to solve optimal control problems. In addition, the model proposed in [18] investigates both purely quantum and hybrid PINNs, highlighting parameter efficiency but relying on limited PQC networks and addressing primarily simple PDEs. Recently proposed HQPINN [19], inspired by the Fourier neural operator (FNO) [65], introduces a parallel hybrid architecture to separately model harmonic and non-harmonic features before combining them for the output. The present work addresses general PDE solving with multiple quantum circuit configurations and explicit emphasis on parameter efficiency.
@@ 3.1 Physics Informed Neural Network (PINNs)
In the present work, we consider a general partial differential equation of the form:
$\mathcal{D}[u(\mathbf{x});\alpha]$ $=f(\mathbf{x}),\quad\mathbf{x}\in\Omega,$ (1)
$\mathcal{B}_{k}[u(\mathbf{x})]$ $=g_{k}(\mathbf{x}),\quad\mathbf{x}\in\Gamma_{k}\subset\partial\Omega,\quad k=1,2,\ldots,n_{b}$
where, $\mathcal{D}$, $\mathcal{B}_{k}$ denote arbitrary differential and boundary/initial operators, respectively. The term $f(\mathbf{x})$ refers to the source term or forcing function in the PDE. The functions $g_{k}(\mathbf{x})$ refer to a set of boundary/initial functions, $\alpha$ is a set of parameter vector or coefficient and $u(\mathbf{x})$ is the solution of the differential equation given input $\mathbf{x}$ in the solution domain $\Omega$ with boundary $\partial\Omega$.
PINNs [66] provide a framework to embed governing physical laws directly into neural network architectures. The PINN loss function combines terms representing governing equations and initial/boundary conditions designed to minimize deviations from physical laws. More formally, the PINN loss function can be defined as:
$\begin{split}\mathcal{L}(\theta)=&\arg\min_{\theta}\sum_{k=1}^{n}\lambda_{k}\mathcal{L}_{k}(\theta)\\ =&\arg\min_{\theta}\left(\lambda_{1}\mathcal{L}_{1}(D[u_{\theta}(\mathbf{x});\mathbf{\alpha}]-f(\mathbf{x}))+\sum_{k=2}^{n_{b}}\lambda_{k}\mathcal{L}_{k}(B[u_{\theta}(\mathbf{x})]-g_{k}(\mathbf{x}))\right),\end{split}$ (2)
where, $n_{b}$ denotes the total number of loss terms, $\mathcal{L}_{k}$ are the individual loss functions for each constraint with weighting coefficients $\lambda_{k}$, and $\theta$ represents the trainable parameters.
While the initial and boundary conditions guarantee a unique solution by satisfies predefined constraints at specific spatial or temporal locations, the physics loss acts as a regularizer, guiding the network to learn physically consistent solutions and preventing overfitting to sparse or noisy data. The balance between these loss components is crucial for achieving accurate solutions, as improper weighting can lead to underfitting the initial/boundary conditions or failing to capture the system dynamics correctly. In the present work adopts soft imposition of the boundary and initial conditions using fixed weights with the best empirical results. However, various weighting methods have been proposed in the literature using different criteria, such as gradient information [67, 68, 69], loss-term history statistics [70, 71], training dataset information [72, 73, 74], or hard imposition of initial and boundary conditions [75, 76],to name a few.
@@ 3.2 Continuous Variable (CV) QNN
In photonic quantum computing, CV computation addresses quantum states with continuous degrees of freedom, such as the position ($x$) and momentum ($p$) quadratures of the electromagnetic field. Within this framework, qumodes represent the fundamental units of the CV quantum systems, analogous to qubits in discrete-variable systems. Unlike qubits, qumodes encode information in the continuous degrees of freedom of electromagnetic field modes, rather than discrete two-level states. The computational space is, in principle, unbounded, involving infinite-dimensional Hilbert spaces with continuous spectra. To make these simulations computationally feasible, a cutoff dimension is employed to limit the number of basis states used for approximation. The computational state space is defined as $r=m^{n}$, with $n$ qumodes and cutoff dimension $m$. For example, with three qumodes and a cutoff dimension of 2, the state space is 8-dimensional. Ideally, higher cutoff dimensions (up to a certain threshold) yield more precise results but at the cost of increased computational complexity.
@@
where, $M$ represents the combined symplectic matrix of the interferometers and squeezing gates, $x$ is the input vector in the phase space, $b$ is the displacement vector with nonlinear transformation $\Phi$.
In addition to the building blocks above, CV-Circuit QNNs offer flexibility in measurement schemes and parameterization strategies. Measurement can be performed in either the number basis, which projects the state onto discrete photon-number states, or the quadrature basis, which yields continuous-valued outcomes corresponding to position or momentum observables. The parameterization of unitary operations may either be full (including both magnitude and phase) or restricted to magnitude-only (by setting phase parameters $\phi$ to zero), which can improve trainability and mitigate optimization challenges at the cost of reduced expressivity.
Although such a model exhibits intriguing theoretical properties, their practical implementation in real-world problem solving seems to suffer from fundamental stability issues. This is mainly due to their inherent sensitivity and reliance on continuous degrees of freedom, which complicate optimization compared to DV-based approaches (discussed in detail in section 3.3). This sensitivity stems from several factors, including the precision required for state preparation, noise from decoherence, and the accumulation of errors in variational quantum circuits. Furthermore, the transformations involved in CV quantum circuits are prone to numerical instabilities, particularly when optimized using gradient-based methods. Although foundational studies about the barren plateau problem, such as [77, 78, 79], primarily address DV circuit systems, the analysis indicates that similar issues can arise in CV quantum systems. In particular, the infinite-dimensional Hilbert space and the continuous nature of parameterization can exacerbate the barren-plateau effects.
@@ 3.3 Discrete Variable (DV) QNN
DV-Circuit quantum models do not inherently provide affine transformations as a core feature like CV-based QNNs discussed in 3.2. However, they can achieve affine-like behaviour through parameterized single-qubit rotations, controlled operations, and phase shifts for amplitude adjustments. Therefore, the affine transform of equation (3) can be reinterpreted to align with their architecture such that:
@@
    - $U_{1},U_{2}$: A unitary operator that can be represented by parameterized single-qubit rotations, such as $RX(\theta),RY(\theta)$, or $RZ(\theta)$.
    - $S$: A sequence of controlled gates (e.g., CNOT or CZ) responsible for introducing entanglement, which enables the representation of correlated variables and interactions.